# Quantifying and Optimizing Global Faithfulness in Persona-driven Role-playing

### v1

@@ 1 Introduction
Role-playing (Li et al., 2023; Yan et al., 2023; Bianchi et al., 2024; Yu et al., 2024; Tao et al., 2024) is a newborn and trending natural language processing field, emerging from the proficiency of large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023a, b; Mesnard et al., 2024) in human interaction. Role-playing customized AI characters, which are useful for providing emotional value (Zhang et al., 2024), developing video games (Hu et al., 2024), or even realizing the metaverse (Zhou, 2023; Yue et al., 2024). Persona-driven role-playing (PRP) (Shao et al., 2023; Xu et al., 2024) uses only persona statements to efficiently build the AI character without dialogues or scripts, which is extremely useful for real-world applications as few characters have sufficient or accessible dialogues for training.
@@
Figure 1: A presentation of the alignment between the APC definition and the human’s view on PRP faithfulness constraints.
Our key insights are 1) the statement-to-response constraint depends on query-statement relevance and 2) the statement-to-response constraint can be formalized as statement-to-response natural language inference (NLI). (Bowman et al., 2015) The constraint becomes active when the query is relevant to the persona statement, constraining the response to be entailed by the persona statement. The constraint becomes passive when the query is irrelevant to the persona statement, reducing the constraint to only not being contradicted by the persona statement. We present a possible PRP instance in Figure 1 to show how our definition is consistent with human’s view on PRP faithfulness. As $q_{1}$ is relevant to $s$, $s$ becomes active and constrains the character “Alice” to incorporate the information in $s$ to her response. For irrelevant $q_{2}$, the constraint of $s$ becomes passive and is relaxed to only not incorporating information contradicting $s$. The discrimination of this Active-Passive-Constraint (APC) is also consistent with human intuition, establishing itself as a solid criterion for PRP faithfulness evaluation.
@@
Furthermore, we discover the APC score to be a reliable reward for direct preference optimization (DPO) (Rafailov et al., 2023) to strengthen the faithfulness of PRP methods. We use APC and human evaluation to verify the effectiveness of DPO, which benefits the satisfaction of both active and negative constraints. We extend the experiments for evaluation and DPO above to complicated famous figures with $77\sim 599$ persona statements, which further verify the conclusions reached in our experiments.
Finally, we launch case studies toward a specific analysis of the insights obtained by APC score-based evaluation and the benefit gained from APC-based DPO. We also showcase how we can explain the detected constraint violation by tracing back and strengthening extra constraints like protective experience by persona statements.
  • We propose the first formal definition of AI character’s global faithfulness and formulate it as a constraint satisfaction problem.
  • We further provide a quantified formulation, APC score, for the constraint, which is human-consistent and the first quantified evaluation for AI characters.
@@
  • We find APC-based DPO to be one of the most competitive techniques to improve the global faithfulness of AI characters.
@@ 2 Related Works
With the emergence of the high capability of LLMs in interaction with humans, role-playing AI has attracted lots of attention from both academia (Shanahan et al., 2023) and industry [^2] https://character.ai/. The difference between role-playing and normal agents is the demand of following a constant persona. The main aim of role-playing includes personalizing the agent for the user preference (Jang et al., 2023) and bringing virtual characters to the real world (Li et al., 2023; Tao et al., 2024). Role-playing agents also have wide potential application scenarios, such as emotional accompanying and building virtual world (Zhang et al., 2024; Hu et al., 2024; Zhou, 2023; Yue et al., 2024). A straightforward implementation for role-playing is fine-tuning LLMs on the dialogues of the characters (Li et al., 2023), which is limited in broad application since rare characters have sufficient accessible dialogue data for fully mastering the character persona.
Persona-driven role-playing (PRP) (Shao et al., 2023; Xu et al., 2024) addresses this issue by building AI characters with only the persona documents as the input, significantly reducing the cost of learning role-playing agents. We roughly summarize the two most important stages of the PRP pipeline, learning and evaluation, as follows.
@@
is a crucial aspect of PRP systems. Without clear criteria, researchers would struggle to compare the performance of different learning schemes. Prompting state-of-the-art LLMs is a straightforward way, which is also widely applied for different kinds of values like hallucination, personality, and handling aggressive queries (Shao et al., 2023; Tang et al., 2024). Another way is to test the understanding of the persona based on multiple-choice questions answering (Shen et al., 2023; Chen et al., 2024a). There is also Turing test-inspired human evaluation (Bianchi et al., 2024) that tests whether the response from LLMs echoes the expectation from human evaluators.
Unfortunately, these evaluation methods are either vague or indirect. Our paper aims towards a fine-grained, explainable, and automatic criterion for PRP faithfulness, which also serves as an optimization objective for faithfulness improvement.
@@ 3.1 Persona-driven Role-Playing
A persona-driven role-playing (PRP) agent (AI character) is defined as a function $f(\cdot)$ that $r\sim f(q|S)$, which generates a response $r$ to a query $q$ (including the history in multi-turn interactions), referring to the list of persona statements $S=[s_{1},s_{2},\cdots,s_{|S|}]$. Ideally, each persona statement should be atomic, indicating it includes only one fact (attribute, experience, etc.) about the character. Existing PRP agents are mostly based on LLMs, denoted as $f_{\mathrm{LLM}}(\cdot)$, which takes a prompt as the input and outputs a response.
@@ 3.3 Experience Upload
Experience upload (EU) is another way to build an AI agent without persona statements inside the input prompt. For each persona statement $s$, EU prompts the LLM to generate $(q,r)$ pairs that $q$ is generally relevant to $s$ and $r$ is faithful to $s$. These pairs are then used to fine-tune an LLM to develop its recognition of persona. On the role-playing stage, the LLM only takes the query as input, $r\sim f_{\mathrm{LLM}}(q)$.
@@ 4 Active-Passive-Constraint
While several potential methods exist for building a PRP agent, there isn’t a clear criterion to accurately assess how faithfully the response $r$ reflects the intended output. Existing criteria generally ask humans or LLMs to score the response by their memory of the character’s persona, which is coarse-grained and vague. Our metric, the Active-Passive-Constraint (APC) score aims to establish a fine-grained and explainable criterion for PRP faithfulness.
@@ 4.1 Definition and Formulation
We first recall the high-level idea of APC mentioned in the introduction that we aim to formulate faithful PRP as a constraint satisfaction problem (CSP). For each persona statement $s$ as constraint, the condition to satisfy it depends on whether it is relevant to the query $q$ (active) or not (passive). We introduce a Boolean function $g(\cdot)$ to represent this status, $g(s,q)$ returns $1$ when $s$, $q$ are relevant and returns 00 for irrelevance. When the constraint is active ($g(s,q)=1$), the response $r$ is constrained to be entailed by $s$, denoted as $s\models r$ (MacCartney & Manning, 2014). When the constraint is passive ($g(s,q)=0$) in natural language inference (NLI), the constraint for $r$ is released to only not being contradicted by $s$ in NLI, denoted as $s\not\models\neg r$ (MacCartney & Manning, 2014). As the semantics of $r$ is affected by $q$, we also introduce $q$ as a condition for the persona statement-to-response NLI, resulting in the following APC constraint for each person statement $s$.
$\textrm{APC}(q,r|s)=(g(s,q)\land(s\models r|q))\lor(\neg g(s,q)\land(s\not\models\neg r|q))$
@@
$\textrm{APC}(q,r|S)=\land_{i=1}^{|S|}\textrm{APC}(q,r|s_{i})=\land_{i=1}^{|S|}\left[(g(s_{i},q)\land(s_{i}\models r|q))\lor(\neg g(s_{i},q)\land(s_{i}\not\models\neg r|q))\right]$
@@ 4.2 Mathematical Quantification
While APC directly discriminates whether a response $r$ is faithful to all persona statements $S$, its strictness hinders its application to PRP agent comparison. Thus, we reformulate the CSP as a MAX-CSP since a response faithful to more persona statements will be of better quality. The metric, APC score ($V_{\mathrm{APC}}(\cdot)$) counts the number of constraints satisfied by the response. To further fine-granularize the metric, we introduce a probabilistic system $P_{\mathrm{APC}}(\cdot)$, which evaluates the probability of each constraint being satisfied.
$V_{\mathrm{APC}}(q,r|S)=\#_{i=1,\cdots,|S|}[\textrm{APC}(q,r|s_{i})]=\sum_{i=1}^{|S|}P_{\mathrm{APC}}(q,r|s_{i})$
@@
$P_{\mathrm{APC}}(q,r|s_{i})=(P_{g}(s_{i},q)P_{h}(s_{i}\models r|q))+(1-P_{g}(s_{i},q))P_{h}(s_{i}\not\models\neg r|q)$
@@
$V_{\mathrm{APC}}(q,r|S)=\sum_{i=1}^{|S|}[(P_{g}(s_{i},q)P_{h}(s_{i}\models r|q))+(1-P_{g}(s_{i},q))P_{h}(s_{i}\not\models\neg r|q)]$
@@
$\Delta V_{\mathrm{APC}}(q,r|S)=V_{\mathrm{APC}}(q,r|S)-\sum_{i=1}^{|S|}(1-P_{g}(s_{i},q))$
@@ 4.3 Weakness of PRP Methods from APC’s View
Figure 2: An overview of different PRP methods.
From APC’s view of PRP faithfulness, we can gain high-level insights into the weakness of existing PRP methods depicted in Figure 2.
@@
Understanding the weaknesses of the methods above is beneficial for us to propose new methods to optimize the global faithfulness of PRP.
@@ 4.4 APC-based Direct Preference Optimization
Our APC score can also act as a reward model for direct preference optimization (DPO), whose initial formulation is presented as follows.
$\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta},\pi_{\mathrm{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}\right)\right]$
@@
$\pi_{\mathrm{ref}}(r|g(s,q))=\pi_{(a)}(r|q,s_{i})=P_{h}(q\models r|s_{i});\pi_{\mathrm{ref}}(r|\neg g(s,q))=\pi_{(p)}(r|q,s_{i})=P_{h}(q\not\models\neg r|s_{i}).$
@@
$\mathcal{L}_{\mathrm{APC}}(\pi_{\theta},\pi_{(a)},\pi_{(p)})=\sum_{i=1}^{|S|}P_{g}(s_{i},q)\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta},\pi_{({a})})+(1-P_{g}(s_{i},q))\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta},\pi_{({p})})$
@@ 5.1 Implementation Details
We follow Shao et al. (2023) to evaluate PRP agents by interview but take the APC score as the metric for responses. We implement the APC score criterion by symbolically distilling from the state-of-the-art LLM, GPT-4 (OpenAI, 2023) and report the regularized $\mathbf{\Delta}$APC score. For statement-query relevance and statement-to-response NLI, we fill in templates with input information shown in the Appendix G and prompt GPT-4 to output the label. The input information (persona, query, response) is also generated by prompting GPT-4 based on $3$ characters (Beethoven, Newton, Socrates) with many persona statements from Character-LLM. We got $8.4K$ data for statement-query relevance and $18.9K$ data for statement-to-response NLI, which are used to fine-tune a state-of-the-art discriminator DeBERTa-V3 ($\sim$300M parameters) (He et al., 2021) for efficiency. We use $80\%$/$20\%$ train/test split and observe a high ($\sim 90\%$) accuracy referencing GPT-4’s labels, which guarantees a high capability of the distilled discriminator. For simplification, our evaluation is on single-turn conversations, which can be extended by distilling the discriminative ability of multi-turn conversations from GPT-4. More details about the distillation can be found in the Appendix B. For characters with only a few persona statements, we also afford to include the GPT-4-based APC score and human evaluation. The human evaluators are asked to memorize these few persona statements and assign quality scores to responses to analyze human alignment. The human evaluator follows a $10$-score scheme with more details in the Appendix D.
@@ 5.2 Compared Methods
We include different PRP methods for evaluation to analyze their advantages and limitations. All methods, except prompting closed-source LLMs, use Gemma (Gemma-1.1-7B-it) as the PRP foundation LLM and low-rank optimization (Hu et al., 2021).
@@
  • APC-based Direct Preference Optimization assigns preference to sampled responses from PRP methods by APC score. The training is retrained to be evaluator-agnostic, which uses a character-specific APC scoring system detailed in Appendix B for fairness. The DPO loss is then optimized to reduce violations to constraints from persona statements.
The setup of hyperparameters of these methods can be found in the Appendix C for reproduction. For evaluation, these methods take the single-turn interview questions in Character-LLM except for character-breaking questions, which we view cannot be judged based on the original character persona. We further discuss how to inject protective persona statements to handle those questions in Section 6.3.

### v2

@@ 1 Introduction
Role-playing (Han et al., 2022; Li et al., 2023; Yan et al., 2023; Bianchi et al., 2024; Yu et al., 2024; Tao et al., 2024) is a newborn and trending natural language processing field, emerging from the proficiency of large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023a, b; Mesnard et al., 2024) in human interaction. Role-playing customized AI characters, which are useful for providing emotional value (Zhang et al., 2024), developing video games (Hu et al., 2024), or even realizing the metaverse (Zhou, 2023; Yue et al., 2024). Persona-driven role-playing (PRP) (Wang et al., 2023a, b; Shao et al., 2023; Xu et al., 2024) uses only persona statements to efficiently build the AI character without dialogues or scripts, which is extremely useful for real-world applications as few characters have sufficient or accessible dialogues for training.
@@
Figure 1: A presentation of the alignment between APC and human’s view on PRP faithfulness.
Our key insights are 1) the statement-to-response constraint depends on query-statement relevance and 2) the statement-to-response constraint can be formalized as statement-to-response natural language inference (NLI). (Bowman et al., 2015) The constraint becomes active when the query is relevant to the persona statement, constraining the response to be entailed by the persona statement. The constraint becomes passive when the query is irrelevant to the persona statement, reducing the constraint to only not being contradicted by the persona statement. We present a possible PRP instance in Figure 1 to show how our definition is consistent with human’s view on PRP faithfulness. As $q_{1}$ is relevant to $s$, $s$ becomes active and constrains the character “Alice” to incorporate the information in $s$ to her response. For irrelevant $q_{2}$, the constraint of $s$ becomes passive and is relaxed to only not incorporating information contradicting $s$.
@@
Furthermore, we discover the APC score to be a reliable reward for direct preference optimization (DPO) (Rafailov et al., 2023) to strengthen the faithfulness of PRP methods. We use APC and human evaluation to verify the effectiveness of DPO, which benefits the satisfaction of both active and negative constraints. We extend the experiments for evaluation and DPO above to complicated famous figures with $77\sim 599$ persona statements, further verifying the reached conclusions.
Finally, we launch case studies toward a specific analysis of the insights obtained by APC score-based evaluation and the benefit gained from APC-based DPO. We also showcase how we can explain the detected constraint violation by tracing back and strengthening extra constraints like protective experience by persona statements. Our contribution is three-fold,
  • We propose the first formal definition of AI character’s global faithfulness and formulate it as a constraint satisfaction problem. The constraint is further quantified as the APC score, which is human-consistent and the first quantified evaluation for AI characters.
@@
  • We find APC-based DPO to be one of the most competitive techniques to improve the global faithfulness of AI characters and cooperate well with other methods.
@@ 2 Related Works
With the emergence of the high capability of LLMs in interaction with humans, role-playing AI has attracted lots of attention from both academia (Shanahan et al., 2023) and industry [^2] https://character.ai/. The difference between role-playing and normal agents is the demand of following a constant persona. The main aim of role-playing includes personalizing the agent for the user preference (Jang et al., 2023) and bringing virtual characters to the real world (Li et al., 2023; Tao et al., 2024). Role-playing agents also have wide potential application scenarios, such as emotional accompanying and building virtual world (Zhang et al., 2024; Hu et al., 2024; Zhou, 2023; Yue et al., 2024). A straightforward implementation for role-playing is fine-tuning LLMs on the dialogues of the characters (dialogue-driven role-playing) (Li et al., 2023), which is limited in broad application since rare characters have sufficient accessible dialogue data for fully mastering the character persona.
Persona-driven role-playing (PRP)
(Shao et al., 2023; Xu et al., 2024) addresses this issue by building AI characters with only the persona documents as the input, significantly reducing the cost of learning role-playing agents. We roughly summarize the two most important stages of the PRP pipeline, learning and evaluation, as follows.
@@
is a crucial aspect of PRP systems. Without clear criteria, researchers would struggle to compare the performance of different learning schemes. Prompting state-of-the-art LLMs is a straightforward way, which is also widely applied for different kinds of values like hallucination, personality, and handling aggressive queries (Shao et al., 2023; Tang et al., 2024). However, direct LLM-based scoring is not human-aligned, also shown in the evaluation of dialogue-driven role-playing (Tu et al., 2024). Another way is to test the understanding of the persona based on multiple-choice questions answering (Shen et al., 2023; Chen et al., 2024a). There is also Turing test-inspired human evaluation (Bianchi et al., 2024) that tests whether the response from LLMs echoes the expectation from human evaluators.
Unfortunately, these evaluation methods for PRP are either vague or indirect. Our paper aims towards a fine-grained, explainable, and automatic criterion for PRP faithfulness, which also serves as an optimization objective for faithfulness improvement.
@@ 3.1 Persona-driven Role-Playing
A persona-driven role-playing (PRP) agent (AI character) is defined as a function $f(\cdot)$ that $r\sim f(q|S)$, which generates a response $r$ to a query $q$ (including the history in multi-turn interactions), referring to persona statements $S=[s_{1},s_{2},\cdots,s_{|S|}]$. Ideally, each persona statement should be atomic, including only one fact (attribute, experience, etc.) about the character. Existing PRP agents are mostly based on LLMs, denoted as $f_{\mathrm{LLM}}(\cdot)$, taking a prompt as the input and outputs a response.
@@ 3.3 Experience Upload
Experience upload (EU) (Shao et al., 2023) is another way to build an AI agent without persona statements inside the input prompt. For each persona statement $s$, EU prompts the LLM to generate $(q,r)$ pairs that $q$ is generally relevant to $s$ and $r$ is faithful to $s$. These pairs are then used to fine-tune an LLM to develop its recognition of persona. On the role-playing stage, the LLM only takes the query as input, $r\sim f_{\mathrm{LLM}}(q)$.
@@ 4.1 Definition and Formulation
We first recall the high-level idea of APC mentioned in the introduction that we aim to formulate faithful PRP as a constraint satisfaction problem (CSP). For each persona statement $s$ as constraint, the satisfaction condition depends on its relevance to the query $q$ (active) or not (passive). We introduce a Boolean function $g(\cdot)$ to represent this status, $g(s,q)$ returns $1$ when $s$, $q$ are relevant and returns 00 for irrelevance. When the constraint is active ($g(s,q)=1$), the response $r$ is constrained to be entailed by $s$, denoted as $s\models r$ (MacCartney & Manning, 2014). When the constraint is passive ($g(s,q)=0$) in natural language inference (NLI), the constraint for $r$ is released to only not being contradicted by $s$, denoted as $s\not\models\neg r$. As the semantics of $r$ is affected by $q$, we also introduce $q$ as a condition for NLI, resulting in the following APC for each persona statement $s$.
$\small\textrm{APC}(q,r|s)=(g(s,q)\land(s\models r|q))\lor(\neg g(s,q)\land(s\not\models\neg r|q))$ (1)
@@
$\small\textrm{APC}(q,r|S)=\land_{i=1}^{|S|}\textrm{APC}(q,r|s_{i})=\land_{i=1}^{|S|}\left[(g(s_{i},q)\land(s_{i}\models r|q))\lor(\neg g(s_{i},q)\land(s_{i}\not\models\neg r|q))\right]$ (2)
@@ 4.2 Mathematical Quantification
While APC directly discriminates whether a response $r$ is faithful to all persona statements $S$, its strictness hinders its application to PRP agent comparison. Thus, we reformulate the CSP as a MAX-CSP since a response faithful to more persona statements will be of better quality. The metric, APC score ($V_{\mathrm{APC}}(\cdot)$) counts the number of constraints satisfied by the response. To further fine-granularize the metric, we introduce $P_{\mathrm{APC}}(\cdot)$ evaluating the probability of each constraint being satisfied.
$\small V_{\mathrm{APC}}(q,r|S)=\#_{i=1,\cdots,|S|}[\textrm{APC}(q,r|s_{i})]=\sum_{i=1}^{|S|}P_{\mathrm{APC}}(q,r|s_{i})$ (3)
@@
$\small P_{\mathrm{APC}}(q,r|s_{i})=(P_{g}(s_{i},q)P_{h}(s_{i}\models r|q))+(1-P_{g}(s_{i},q))P_{h}(s_{i}\not\models\neg r|q)$ (4)
@@
$\small V_{\mathrm{APC}}(q,r|S)=\sum_{i=1}^{|S|}[(P_{g}(s_{i},q)P_{h}(s_{i}\models r|q))+(1-P_{g}(s_{i},q))P_{h}(s_{i}\not\models\neg r|q)]$ (5)
@@
$\Delta V_{\mathrm{APC}}(q,r|S)=V_{\mathrm{APC}}(q,r|S)-\sum_{i=1}^{|S|}(1-P_{g}(s_{i},q))$ (6)
@@ 4.3 Weakness of PRP Methods from APC’s View
From APC’s view of PRP faithfulness, we can gain insights into the weakness of PRP techniques.
@@ 4.4 APC-based Direct Preference Optimization
Our APC score also acts as a reward for direct preference optimization (DPO) (Rafailov et al., 2023), whose initial formulation is presented as follows.
$\small\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta},\pi_{\mathrm{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}\right)\right]$ (7)
@@
$\small\pi_{\mathrm{ref}}(r|g(s,q))=\pi_{(a)}(r|q,s_{i})=P_{h}(q\models r|s_{i});\pi_{\mathrm{ref}}(r|\neg g(s,q))=\pi_{(p)}(r|q,s_{i})=P_{h}(q\not\models\neg r|s_{i}).$ (8)
@@
$\small\mathcal{L}_{\mathrm{APC}}(\pi_{\theta},\pi_{(a)},\pi_{(p)})=\sum_{i=1}^{|S|}P_{g}(s_{i},q)\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta},\pi_{({a})})+(1-P_{g}(s_{i},q))\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta},\pi_{({p})})$ (9)
@@ 5.1 Implementation Details
We follow Shao et al. (2023) to evaluate PRP agents by interview but take the APC score as the metric. We implement the APC score criterion by symbolically distilling from the state-of-the-art LLM, GPT-4 (OpenAI, 2023) and report the regularized $\mathbf{\Delta}$APC score. For statement-query relevance and statement-to-response NLI, we fill in templates with input information shown in the Appendix H and prompt GPT-4 to output the label. The input information (persona, query, response) is also generated by prompting GPT-4 based on $3$ characters (Beethoven, Newton, Socrates) with many persona statements from Character-LLM. We got $8.4K$ data for statement-query relevance and $18.9K$ data for statement-to-response NLI, which are used to fine-tune a state-of-the-art discriminator DeBERTa-V3 ($\sim$300M parameters) (He et al., 2021) for efficiency. We use $80\%$/$20\%$ train/test split and observe a high ($\sim 90\%$) accuracy referencing GPT-4’s labels, which guarantees a high capability of the distilled discriminator. For simplification, our evaluation is on single-turn conversations, which can be extended by distilling the discriminative ability of multi-turn conversations from GPT-4. More details about the distillation can be found in the Appendix C. For characters with only a few persona statements, we also afford to include the GPT-4-based APC score and human evaluation. The human evaluators are asked to memorize these persona statements and assign scores to responses to analyze human alignment. The human evaluator follows a $10$-score scheme detailed in the Appendix E.
@@ 5.2 Compared Methods
We include different PRP methods for evaluation to analyze their advantages and limitations. All methods, except prompting closed-source LLMs, use Gemma (Gemma-1.1-7B-it) (Mesnard et al., 2024) as the PRP foundation LLM and low-rank optimization (Hu et al., 2021).
@@
  • APC-based Direct Preference Optimization assigns preference to sampled responses from PRP methods by APC score. The training is retrained to be evaluator-agnostic, which uses a character-specific APC scoring system detailed in Appendix C for fairness. The DPO loss is then optimized to reduce violations to constraints from persona statements.
The setup of hyperparameters can be found in the Appendix D for reproduction. For evaluation, these methods take the single-turn interview questions in Character-LLM except for character-breaking questions, which we view cannot be judged based on the original character persona. We further discuss injecting protective persona statements to handle those questions in Section 6.3.